OpenStudy (anonymous):

What is the exact volume of a cylinder that has a height of 1.8 ft and a radius of 6.1 ft?

OpenStudy (anonymous):

What's the problem to just do it yourself? Maybe you'd learn something that way. Area = length x width.

OpenStudy (anonymous):

\(\sf V = \pi r^2 h\) Just ignore pi and plug in the radius and height. \(\sf V = \pi (6.1)^2(1.8)\) Simplify

OpenStudy (anonymous):

210.31?

OpenStudy (anonymous):

I said ignore pi..just simplify (6.1^2)(1.8)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

66.978

OpenStudy (anonymous):

Yep, that's our answer. \(\sf 66.978\pi\)